A tuning fork vibrating at a frequency of 512 Hz is held on top of a jar filled with water and fitted with a tap at the bottom. If the jar is 60 cm high and the speed of sound is 350 m/s, describe what occurs as the jar slowly emptied.
1
Expert's answer
2020-05-25T11:18:17-0400

Calculate the wavelength:


"\lambda=\frac{v}{f}=0.684\text{ m}."

As the jar is emptied, the first resonance will occur at


"L_1=\frac{1}{4}\lambda=0.171\text{ m},"

the second will occur at


"L_1=\frac{3}{4}\lambda=0.513\text{ m}."